Please help with these questions

Question 6 0.4 pts When a hurricane rips through Florida, the price of oranges rises because O demand curve shifts to the right. supply curve shifts to the left. O demand curve shifts to the left. Osupply curve shifts to the right. supply and demand curves both shift to the left. Question 7 0.4 pts show(s) how resources and final goods and services flow through the economy Money Circular flow OAllocation OTrade-offs Question 8 0.4 pts As a new firm in the apple-picking business, Nicolette has considered adding an economist to her management team. This economist would be unable to help her managerial team determine O the lowest cost way of distributing apples. O why people eat apples. O the lowest cost way of picking apples. O how many apples consumers will purchase at different prices O the effect government regulations would have on the price of apples.
Answer #1

6. Supply curve shifts to the left.

Hurricane destroys the crop and thus supply decreases which shifts supply curve leftwards.

7. Circular flow

Circular flow depicts the flow of goods and services among various sectors like households, firms, government, rest of the world.

8. why people eat apples

Economists can tell you about effect of government policy on apple consumption, consumption of apples when there is price discrimination, but cannot tell you why people eat apples.

9. firms leaving the market

When firms exit the market then supply of goods decreases which shifts supply curve leftwards.

10. Cross price Ed = % change in quantity demanded of cupcakes / % change in price of muffin

= [(65 - 55)/(65 + 55)/2 x 100]/[(2 - 1)/(2 + 1)/2 x 100]

= (10/60 x 100)/(1/3 x 100) = 16.67/33.33 = 0.50